40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 15568     Entry time: Thu Sep 10 15:56:08 2020
Author: Koji 
Type: Update 
Category: General 
Subject: HEPA & Particle Level Status 

15:30
- PSL HEPA was running at 33% and is now at 100%
- South End HEPA was not on and is now running
- Yarm Portable HEPA was not running and is now running at max speed: the power was taken beneath the ITMY table. It is better to unplug it when one uses the IFO.
- Yend Portable HEPA was not running and is now running (presumably) at max speed

Particle Levels: (Not sure about the unit. The convention here is to multiply x10 of the reading)

Before running the HEPAs at their maximum
9/10/2020 15:30 / 0.3um 292180 / 0.5um 14420

(cf 9/5/2020 / 0.3um 94990 / 0.5um 6210)
==>
After running the HEPAs at their maximum
The number gradually went down and now became constant at about half of the initial values
9/10/2020 19:30 / 0.3um 124400 / 0.5um 7410

ELOG V3.1.3-